# Total trades, intersection matrices and Specht modules

### v1

@@ Abstract.
Trades are important objects in combinatorial design theory that may be realized as certain elements of kernels of inclusion matrices. Total trades were introduced recently by Ghorbani, Kamali and Khosravshahi, who showed that over a field of characteristic zero the vector space of trades decomposes into a direct sum of spaces of total trades. In this paper, we show that the vector space spanned by the permutations of a total trade is an irreducible representation of the symmetric group. As a corollary, the previous decomposition theorem is recovered. Also, a basis is obtained for the module of total trades in the spirit of Specht polynomials. More generally, in the second part of the paper we consider intersection matrices and determine the irreducible decompositions of their images. This generalizes previously known results concerning ranks of special cases.
@@ 1. Introduction
Throughout this paper, we work over a field $\mathbb{K}$ of characteristic zero. Let $t,k,n$ be integers such that $0\leq t<k<n$ and let $X$ be the set $\{1,2,\dots,n\}$. Suppose that $T_{+}$ and $T_{-}$ are two disjoint collections of $k$-subsets of $X$ with the property that the number of occurrences of every $t$-subset of $X$ in $T_{+}$ and $T_{-}$ are the same. Then $T=(T_{+},T_{-})$ is called a $t$-$(n,k)$ trade.
The inclusion matrix $W_{t,k}$ is a matrix with rows and columns indexed by $t$-subsets and $k$-subsets of $X$ respectively and the $(A,B)$ entry is equal to 1 is $A\subseteq B$ and is equal to 0 otherwise. Trades are null vectors of $W_{t,k}$. Among trades, the minimal trades play a crucial role in design theory. A minimal $t$-$(n,k)$ trade may be represented in the form
@@
where $x_{i},y_{i}\in X$ are distinct and multiplication is to be interpreted as union os sets (see Section 2.1 below). The symmetric group $\mathfrak{S}_{n}$ acts naturally on $X$ and thus on minimal trades. Their importance is highlighted by the classical result of Graver and Jurkat [10] that the vector space of $t$-$(n,k)$ trades is generated by the permutations of any minimal $t$-$(n,k)$ trade if $t+k\leq n$.
Recently Ghorbani, Kamali and Khosravshahi [7] introduced the notion of total trade. A $t$-$(n,k)$ total trade may be represented in the form
@@
where the sum ranges over all $(k-t-1)$-subsets $\{x_{i_{t+2}}\cdots x_{i_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t+1},y_{t+1}\}$. Let us denote by $\mathfrak{T}_{t,k,n}$ the set of $t$-$(n,k)$ total trades and by $\langle\mathfrak{T}_{t,k,n}\rangle$ the vector space generated by $\mathfrak{T}_{t,k,n}$. The following result was shown in [7].
@@
If $2k-1\leq n$, then the vector space of $t$-$(n,k)$ trades decomposes as the direct sum $\langle\mathfrak{T}_{t,k,n}\rangle\oplus\cdots\oplus\langle\mathfrak{T}_{k-1,k,n}\rangle$.
@@
We recall that the irreducible modules of the symmetric group $\mathfrak{S}_{n}$ are naturally parametrized by partitions of $n$. For a partition $\lambda$ of $n$ we denote by $S^{\lambda}$ the corresponding irreducible module. This is called a Specht module. In our first main result we prove that $\langle\mathfrak{T}_{i,k,n}\rangle$ is an irreducible module of the symmetric group $\mathfrak{S}_{n}$ (see Theorem 3.4). The precise statement is the following.
@@
Let $t<k$ and $t+k\leq n$. Then as $\mathfrak{S}_{n}$-modules the subspace $\langle\mathfrak{T}_{t,k,n}\rangle$ of $M_{k}$ spanned by the set $\mathfrak{T}_{t,k,n}$ of $t$-$(n,k)$ total trades is isomorphic to the Specht module $S^{(n-t-1,t+1)}$.
From this we obtain a basis of the $\mathbb{Z}$-module spanned by $\mathfrak{T}_{i,k,n}$ (Corollary 3.5). We also recover the result of Graver and Jurcat mentioned above. Our proofs are independent of [7] and we recover the decomposition in Theorem 1.1 as the irreducible decomposition of the kernel of the inclusion matrix $W_{t,k}$.
Suppose $l$ is an integer such that $0\leq l\leq t$. The intersection matrix $U_{t,k,l}$ is a matrix with rows and columns indexed by $t$-subsets and $k$-subsets of $X$ respectively and the $(A,B)$ entry is equal to 1 is $|A\cap B|=l$ and is equal to 00 otherwise. These generalize the inclusion matrices since $U_{t,k,t}=W_{t,k}$. Such matrices have been studied in different contexts in combinatorics such as design theory and association schemes. Ghareghani, Ghorbani and Mohammad-Noori [6] studied these matrices and found an expression for their rank. The matrices $U_{t,k,l}$ as $l$ varies from 00 to $t$ correspond to basis elements of the homomorphism space $\operatorname{Hom}_{\mathfrak{S}_{n}}(M_{k},M_{t})$ between the permutation modules $M_{k}$ and $M_{t}$ of the symmetric group ${\mathfrak{S}_{n}}$. We generalize the previous result on ranks by considering any linear operator in $\operatorname{Hom}_{\mathfrak{S}_{n}}(M_{k},M_{t})$. In order to state the result we need some notation. Suppose $l$ is a nonnegative integer such that $l\leq t$. For $j=0,1,\dots,t$ we define the integers
@@
$A=\bigoplus_{j\in J(t,k,n)}S^{(n-j,j)},$
where where $J(t,k,n)$ is the subset of $\{0,1,\dots,t\}$ consisting of those $j$ such that there exists $l\in\{0,1,\dots,t\}$ satisfying $c_{l}\neq 0$ and $\lambda_{j}(t,k,n;l)\neq 0$. In particular, the rank of the matrix $A$ is equal to
@@
The paper is organized as follows. Section 2 is devoted to background material concerning mainly representations of the symmetric group. In Section 3 we prove our first main result (see Theorem 3.4) and discuss certain corollaries. The second part the paper is independent of the first part but our techniques are similar. Preliminaries on polynomial representations of the general linear group are considered in Section 4. Our second main result (see Theorem 5.3 and Corollary 5.6) is proved in Section 5.
@@ 2.2. Representations of the symmetric group
A partition of $n$ is a sequence $\lambda=(\lambda_{1},\dots,\lambda_{m})$ of nonnegative integers such $\lambda_{1}\geq\dots\geq\lambda_{m}$ and $\lambda_{1}+\cdots+\lambda_{m}=n$. Since the characteristic of $\mathbb{K}$ is zero, the irreducible representations of $\mathfrak{S}_{n}$ over $\mathbb{K}$ are parametrized by partitions $\lambda$ of $n$. We denote by $S^{\lambda}$ the irreducible module corresponding to $\lambda$. This is called a Specht module. Of the various constructions of $S^{\lambda}$ that are available, we will use one that is based on column tabloids. The facts we need may be found in [5, Chapter 7]. In order for this paper to be reasonably self contained, we present the main points below.
@@
We have highlighted the entries in the second column that were transposed.
Let $\lambda=(\lambda_{1},\lambda_{2})$ be a partition of $n$, $U$ a Young tableau of shape $\lambda$, $q$ the corresponding column tabloid in $\tilde{M}^{\lambda}$ and let $11\leq c\leq\lambda_{1}-$ 1. With this data we define an element $g_{U,c}\in\tilde{M}^{\lambda}$. We distinguish two cases.
  • Suppose $c\leq\lambda_{2}$. Let $U_{i}$ be the Young tableaux obtained from $U$ by exchanging the top element of column $c+1$ with the $i$th element of column $c$. Let $q_{i}$ be the column tabloid corresponding to $U_{i}$. Define $g_{U,c}=q-q_{1}-q_{2}$.
  • Suppose $c>\lambda_{2}$. Let $U_{3}$ be the Young tableaux obtained from $U$ by exchanging the two elements in columns $c$ and $c+1$. Let $q_{3}$ be the corresponding column tabloid. Define $g_{U,c}=q-q_{3}$.
@@
Let $Q^{\lambda}$ be the subspace of $\tilde{M}^{\lambda}$ spanned by all the $g_{U,c}$ as $U$ varies over all Young tableaux of shape $\lambda$ and $11\leq c\leq\lambda_{1}-$ 1.
@@ 3. Total trades
In this section we prove that the space generated by the set of $t$-$(n,k)$ total trades is an irreducible module of the symmetric group $\mathfrak{S}_{n}.$
@@ 3.1. An identity
Th identity is clear for $m=0$. Suppose $m\geq 1$. Let $L$ be the left hand side of (3.2) and let $C:=A\cup\{x,y,z\}$. From Lemma 3.2 we have
@@ 3.2. Main result
Let $t<k$ and $t+k\leq n$. Then as $\mathfrak{S}_{n}$-modules the subspace $\langle\mathfrak{T}_{t,k,n}\rangle$ of $M_{k}$ spanned by the set $\mathfrak{T}_{t,k,n}$ of $t$-$(n,k)$ total trades is isomorphic to the Specht module $S^{(n-t-1,t+1)}$.
@@
$q:=$ $x_{1}$cdots $x_{t+1}$ $z_{t+2}$cdots $z_{n-t-1}$ $y_{1}$cdots $y_{t+1}$
to the $t$-$(n,k)$ total trade
@@
where the sum is over all $(k-t-1)$-subsets $\{x_{i_{t+2}},\dots,x_{i_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t+1},y_{t+1}\}$. Since $T$ is skew symmetric with respect to $x_{i},y_{i}$, where $i=1,\dots,t+1$, the map $h$ is well defined. Moreover it easily follows that $h$ is a surjective map of $\mathfrak{S}_{n}$-modules.
@@
  • The sum $S(1):=\sum x_{j_{t+2}}\cdots x_{j_{k}}$ in $T_{1}$ is over all $(k-t-1)$-subsets $\{x_{j_{t+2}},\dots,x_{j_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t},y_{t},z_{t+2},y_{t+1}\}$.
  • the sum $S(2):=\sum x_{j_{t+2}}\cdots x_{j_{k}}$ in $T_{2}$ is over all $(k-t-1)$-subsets $\{x_{l_{t+2}},\dots,x_{l_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t},y_{t},x_{t+1},z_{t+2}\}$.
Let $S:=\sum x_{i_{t+2}}\cdots x_{i_{k}}$ be the sum in $T$. So this sum is over all $(k-t-1)$-subsets $\{x_{i_{t+2}},\dots,x_{i_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t+1},y_{t+1}\}$. With the notation of Lemma 3.3 we have
@@
Since the map $h$ is surjective, the map $\bar{h}$ is surjective. Now since the $\mathfrak{S}_{n}$-module $S^{(n-t-1,t+1)}$ is irreducible, the map $\bar{h}$ is an isomorphism.∎
@@ 3.3. Corollaries
There is interest in obtaining bases of the $\mathbb{Z}$-module spanned by all $t$-$(n,k)$ trades, for example see [10, 8, 14, 15]. The next result gives a basis of the $\mathbb{Z}$-module spanned by the the $t$-$(n,k)$ total trades.
@@
Let $t<k$ and $t+k\leq n$. Let us fix a total order on the set $X$. Then a basis of the vector space $\langle\mathfrak{T}_{t,k,n}\rangle$ is the set of $t$-$(n,k)$ total trades
@@
Moreover, the above $t$-$(n,k)$ total trades form a basis of the $\mathbb{Z}$-module spanned by $\mathfrak{T}_{t,k,n}$.
@@
$x_{1}$cdots $x_{t+1}$ $z_{t+2}$cdots $z_{n-t-1}$
$y_{1}$cdots $y_{t+1}$
to the $t$-$(n,k)$ total trade $T$ induces an isomorphism $\tilde{M}^{(n-t-1,t+1)}/Q^{(n-t-1,t+1)}\to\langle\mathfrak{T}_{t,k,n}\rangle$. The desired result follows from Theorem 3.4.
@@
It was shown in [7, Theorem 11] that the vector space of $t$-$(n,k)$ trades decomposes as the direct sum $\langle\mathfrak{T}_{t,k,n}\rangle\oplus\cdots\oplus\langle\mathfrak{T}_{k-1,k,n}\rangle$ of subspaces of total trades. We recover this result from Theorem 3.4 as the irreducible decomposition of the space of $t$-$(n,k)$ trades, i.e. as the irreducible decomposition of the kernel of the inclusion matrix $W_{t,k}$.
@@
Let $t<k\leq n/2$. Then vector space of $t$-$(n,k)$ trades decomposes as the direct sum $\langle\mathfrak{T}_{t,k,n}\rangle\oplus\cdots\oplus\langle\mathfrak{T}_{k-1,k,n}\rangle$. [^1] In [7, Theorem 11] the result is stated without assumptions on $k,n$. However, in its proof [7, Theorem 9] is used which has such assumptions. These yield $2k-1\leq n$.
@@
From Theorem 3.4 we have $S^{(n-i-1,i+1)}\cong\langle\mathfrak{T}_{i,k,n}\rangle$ for all $i=t,\dots,k-1$. (The hypothesis $k\leq n/2$ is needed in order to apply Theorem 3.4 for $t$ replaced by $k-1$. Also it is needed so that $(n-k,k)$ is a partition.)∎
For a $t$-$(n,k)$ trade $T$ we denote by $\Pi(T)$ the orbit of $T$ under the action of the symmetric group $\mathfrak{S}_{n}$. The following decomposition result was proved in [7, Theorem 2].
@@
where $I\subseteq\{t,\dots,k-1\}$ consists of integers $i$ such that there is an $i$-trade in $\langle\Pi(T)\rangle$ which is not an $(i+1)$-trade. In particular, $\langle\Pi(T)\rangle$ is the whole vector space of $t$-$(n,k)$ trades if and only if $I=\{t,\dots,k-1\}$.
@@
From Corollary 3.6 it follows that the irreducible decomposition of the $\mathfrak{S}_{n}$-module $\langle\Pi(T)\rangle$ is of the form $\bigoplus_{i\in I}\langle\mathfrak{T}_{i,k,n}\rangle$, where $I$ is a nonempty subset of $\{t,\dots,k-1\}.$ For every $i\in I$, the set $\mathfrak{T}_{i,k,n}$ consists of $i$-$(n,k)$ trades. These are not $(i+1)$-$(n,k)$ trades; indeed, by Theorem 3.4 we have $\langle\mathfrak{T}_{i,k,n}\rangle=S^{(n-i-1,i+1)}$ and this Specht module is not contained in the kernel of the map $\psi_{k}^{(k-i-1)}:M_{k}\to M_{i+1}$ according to (3.3) for $i+1$ in place of $t$.∎
A classical result of Granert and Jurkat states that the $\mathbb{Z}$-module spanned by all $t$-$(n,k)$ trades is spanned by the permutations of any nonzero minimal $t$-$(n,k)$ trade. As an immediate corollary of Theorem 3.4 we recover the vector space version of this result.
@@
Let $t<k\leq n/2$. The vector space of $t$-$(n,k)$ trades is spanned by the permutations of any nonzero mimimal $t$-$(n,k)$ trade.
@@
Since every total trade is a sum of minimal $t$-$(n,k)$ trades, we see from Corollary 3.6 that the vector space of $t$-$(n,k)$ trades is spanned by the minimal $t$-$(n,k)$ trades. From this and the observation that the symmetric group $\mathfrak{S}_{n}$ acts transitively on minimal trades the result follows. ∎

### v2

@@ Abstract.
Trades are important objects in combinatorial design theory that may be realized as certain elements of kernels of inclusion matrices. Total trades were introduced recently by Ghorbani, Kamali and Khosravshahi who showed that over a field of characteristic zero the vector space of trades decomposes into a direct sum of spaces of total trades. In this paper, we show that the vector space spanned by the permutations of a total trade is an irreducible representation of the symmetric group. As a corollary, the previous decomposition theorem is recovered. Also, a basis is obtained for the module of total trades in the spirit of Specht polynomials. In the second part of the paper we consider more generally intersection matrices and determine the irreducible decompositions of their images. This generalizes previously known results concerning ranks of special cases.
@@ 1. Introduction
Throughout this paper, we work over a field $\mathbb{K}$ of characteristic zero. Let $t,k,n$ be integers such that $0\leq t<k<n$ and let $X$ be the set $\{1,2,\dots,n\}$. Suppose that $T_{+}$ and $T_{-}$ are two disjoint collections of $k$-subsets of $X$ with the property that the number of occurrences of every $t$-subset of $X$ in $T_{+}$ and $T_{-}$ are the same. Then $T=(T_{+},T_{-})$ is called a tt-$(n,k)$ trade.
The inclusion matrix $W_{t,k}$ is a matrix with rows and columns indexed by $t$-subsets and $k$-subsets of $X$ respectively and the $(A,B)$ entry is equal to 1 if $A\subseteq B$ and is equal to 0 otherwise. Trades are null vectors of $W_{t,k}$. Among trades, the minimal trades play a crucial role in design theory. A minimal tt-$(n,k)$ trade may be represented in the form
@@
where $x_{i},y_{i}\in X$ are distinct and multiplication is to be interpreted as union of sets (see Section 2.1 below). The symmetric group $\mathfrak{S}_{n}$ acts naturally on $X$ and thus on minimal trades. Their importance is highlighted by the classical result of Graver and Jurkat [10] that the vector space of tt-$(n,k)$ trades is generated by the permutations of any minimal tt-$(n,k)$ trade if $t+k\leq n$.
Recently Ghorbani, Kamali and Khosravshahi [7] introduced the notion of total trade. A tt-$(n,k)$ total trade may be represented in the form
@@
where the sum ranges over all (kt1)(k-t-1)-subsets $\{x_{i_{t+2}}\cdots x_{i_{k}}\}$ of
$X\setminus\{x_{1},y_{1},\dots,x_{t+1},y_{t+1}\}$. Let us denote by $\mathfrak{T}_{t,k,n}$ the set of tt-$(n,k)$ total trades and by $\langle\mathfrak{T}_{t,k,n}\rangle$ the vector space generated by $\mathfrak{T}_{t,k,n}$. The following result was shown in [7].
@@
If $2k-1\leq n$, then the vector space of tt-$(n,k)$ trades decomposes as the direct sum $\langle\mathfrak{T}_{t,k,n}\rangle\oplus\cdots\oplus\langle\mathfrak{T}_{k-1,k,n}\rangle$.
@@
We recall that the irreducible modules of the symmetric group $\mathfrak{S}_{n}$ are naturally parameterized by partitions of $n$. For a partition $\lambda$ of $n$ we denote by $S^{\lambda}$ the corresponding irreducible module. This is called a Specht module. In our first main result we prove that $\langle\mathfrak{T}_{i,k,n}\rangle$ is an irreducible module of the symmetric group $\mathfrak{S}_{n}$ (see Theorem 3.4). The precise statement is the following.
@@
Let $t<k$ and $t+k\leq n$. Then as $\mathfrak{S}_{n}$-modules the subspace $\langle\mathfrak{T}_{t,k,n}\rangle$ of $M_{k}$ spanned by the set $\mathfrak{T}_{t,k,n}$ of tt-$(n,k)$ total trades is isomorphic to the Specht module $S^{(n-t-1,t+1)}$.
From this we obtain a basis of the $\mathbb{Z}$-module spanned by $\mathfrak{T}_{i,k,n}$ (Corollary 3.5). We also recover the result of Graver and Jurkat mentioned above. Our proofs are independent of [7] and we recover the decomposition in Theorem 1.1 as the irreducible decomposition of the kernel of the inclusion matrix $W_{t,k}$.
Suppose $l$ is an integer such that $0\leq l\leq t$. The intersection matrix $U_{t,k,l}$ is a matrix with rows and columns indexed by $t$-subsets and $k$-subsets of $X$ respectively and the $(A,B)$ entry is equal to 1 is $|A\cap B|=l$ and is equal to 0 otherwise. These generalize the inclusion matrices since $U_{t,k,t}=W_{t,k}$. Such matrices have been studied in different contexts in combinatorics such as design theory and association schemes. Ghareghani, Ghorbani and Mohammad-Noori [6] studied these matrices and found an expression for their rank. The matrices $U_{t,k,l}$ as $l$ varies from 0 to $t$ correspond to basis elements of the homomorphism space $\operatorname{Hom}_{\mathfrak{S}_{n}}(M_{k},M_{t})$ between the permutation modules $M_{k}$ and $M_{t}$ of the symmetric group $\mathfrak{S}_{n}$}. We generalize the previous result on ranks by considering any linear operator in $\operatorname{Hom}_{\mathfrak{S}_{n}}(M_{k},M_{t})$. In order to state the result we need some notation. Suppose $l$ is a nonnegative integer such that $l\leq t$. For $j=0,1,\dots,t$ we define the integers
@@
$\bigoplus_{j\in J(t,k,n)}S^{(n-j,j)},$
where $J(t,k,n)$ is the subset of $\{0,1,\dots,t\}$ consisting of those $j$ such that there exists $l\in\{0,1,\dots,t\}$ satisfying $c_{l}\neq 0$ and $\lambda_{j}(t,k,n;l)\neq 0$. In particular, the rank of the matrix $A$ is equal to
@@
The paper is organized as follows. Section 2 is devoted to background material concerning mainly representations of the symmetric group. In Section 3 we prove our first main result (see Theorem 3.4) and discuss certain corollaries. The second part of the paper is independent of the first part but our techniques are similar. Preliminaries on polynomial representations of the general linear group are considered in Section 4. Our second main result (see Theorem 5.3 and Corollary 5.6) is proved in Section 5.
@@ 2.2. Representations of the symmetric group
A partition of $n$ is a sequence $\lambda=(\lambda_{1},\dots,\lambda_{m})$ of nonnegative integers such $\lambda_{1}\geq\dots\geq\lambda_{m}$ and $\lambda_{1}+\cdots+\lambda_{m}=n$. Since the characteristic of $\mathbb{K}$ is zero, the irreducible representations of $\mathfrak{S}_{n}$ over $\mathbb{K}$ are parameterized by partitions $\lambda$ of $n$. We denote by $S^{\lambda}$ the irreducible module corresponding to $\lambda$. This is called a Specht module. Of the various constructions of $S^{\lambda}$ that are available, we will use one that is based on column tabloids. The facts we need may be found in [5, Chapter 7]. In order for this paper to be reasonably self contained, we present the main points below.
@@
We have highlighted the entries in the second column that were exchanged.
Let $\lambda=(\lambda_{1},\lambda_{2})$ be a partition of $n$, $U$ a Young tableau of shape $\lambda$, $q$ the corresponding column tabloid in $\tilde{M}^{\lambda}$ and let $1\leq c\leq\lambda_{1}-1$. With this data we define an element $g_{U,c}\in\tilde{M}^{\lambda}$. We distinguish two cases.
  • Suppose $c\leq\lambda_{2}$. Let $U_{i}$ be the Young tableau obtained from $U$ by exchanging the top element of column $c+1$ with the $i$th element of column $c$. Let $q_{i}$ be the column tabloid corresponding to $U_{i}$. Define $g_{U,c}=q-q_{1}-q_{2}$.
  • Suppose $c>\lambda_{2}$. Let $U_{3}$ be the Young tableau obtained from $U$ by exchanging the two elements in columns $c$ and $c+1$. Let $q_{3}$ be the corresponding column tabloid. Define $g_{U,c}=q-q_{3}$.
@@
Let $Q^{\lambda}$ be the subspace of $\tilde{M}^{\lambda}$ spanned by all the $g_{U,c}$ as $U$ varies over all Young tableaux of shape $\lambda$ and $1\leq c\leq\lambda_{1}-1$.
@@ 3. Total trades
In this section we prove that the space generated by the set of tt-$(n,k)$ total trades is an irreducible module of the symmetric group $\mathfrak{S}_{n}.$
@@ 3.1. An identity
The identity is clear for $m=0$. Suppose $m\geq 1$. Let $L$ be the left hand side of (3.2) and let $C:=A\cup\{x,y,z\}$. From Lemma 3.2 we have
@@ 3.2. Main result
Let $t<k$ and $t+k\leq n$. Then as $\mathfrak{S}_{n}$-modules the subspace $\langle\mathfrak{T}_{t,k,n}\rangle$ of $M_{k}$ spanned by the set $\mathfrak{T}_{t,k,n}$ of tt-$(n,k)$ total trades is isomorphic to the Specht module $S^{(n-t-1,t+1)}$.
@@
$q:=$ $x_{1}$ $\cdots$ $x_{t+1}$ $z_{t+2}$ $\cdots$ $z_{n-t-1}$ $y_{1}$ $\cdots$ $y_{t+1}$
to the tt-$(n,k)$ total trade
@@
where the sum is over all (kt1)(k-t-1)-subsets $\{x_{i_{t+2}},\dots,x_{i_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t+1},y_{t+1}\}$. Since $T$ is skew symmetric with respect to $x_{i},y_{i}$, where $i=1,\dots,t+1$, the map $h$ is well defined. Moreover it easily follows that $h$ is a surjective map of $\mathfrak{S}_{n}$-modules.
@@
  • The sum $S(1):=\sum x_{j_{t+2}}\cdots x_{j_{k}}$ in $T_{1}$ is over all (kt1)(k-t-1)-subsets $\{x_{j_{t+2}},\dots,x_{j_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t},y_{t},z_{t+2},y_{t+1}\}$.
  • the sum $S(2):=\sum x_{j_{t+2}}\cdots x_{j_{k}}$ in $T_{2}$ is over all (kt1)(k-t-1)-subsets $\{x_{l_{t+2}},\dots,x_{l_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t},y_{t},x_{t+1},z_{t+2}\}$.
Let $S:=\sum x_{i_{t+2}}\cdots x_{i_{k}}$ be the sum in $T$. So this sum is over all (kt1)(k-t-1)-subsets $\{x_{i_{t+2}},\dots,x_{i_{k}}\}$ of $X\setminus\{x_{1},y_{1},\dots,x_{t+1},y_{t+1}\}$. With the notation of Lemma 3.3 we have
@@
Since the map $h$ is surjective, the map $\bar{h}$ is surjective. Now since the $\mathfrak{S}_{n}$-module $S^{(n-t-1,t+1)}$ is irreducible and $\langle\mathfrak{T}_{t,k,n}\rangle$ is nonzero , the map $\bar{h}$ is an isomorphism.∎
@@ 3.3. Corollaries
There is interest in obtaining bases of the $\mathbb{Z}$-module spanned by all tt-$(n,k)$ trades, for example see [10, 8, 14, 15]. The next result gives a basis of the $\mathbb{Z}$-module spanned by the the tt-$(n,k)$ total trades.
@@
Let $t<k$ and $t+k\leq n$. Let us fix a total order on the set $X$. Then a basis of the vector space $\langle\mathfrak{T}_{t,k,n}\rangle$ is the set of tt-$(n,k)$ total trades
@@
Moreover, the above tt-$(n,k)$ total trades form a basis of the $\mathbb{Z}$-module spanned by $\mathfrak{T}_{t,k,n}$.
@@
$x_{1}$ $\cdots$ $x_{t+1}$ $z_{t+2}$ $\cdots$ $z_{n-t-1}$
$y_{1}$ $\cdots$ $y_{t+1}$
to the tt-$(n,k)$ total trade $T$ induces an isomorphism $\tilde{M}^{(n-t-1,t+1)}/Q^{(n-t-1,t+1)}\to\langle\mathfrak{T}_{t,k,n}\rangle$. The desired result follows from Theorem 3.4.
@@
It was shown in [7, Theorem 11] that the vector space of tt-$(n,k)$ trades decomposes as the direct sum $\langle\mathfrak{T}_{t,k,n}\rangle\oplus\cdots\oplus\langle\mathfrak{T}_{k-1,k,n}\rangle$ of subspaces of total trades. We recover this result from Theorem 3.4 as the irreducible decomposition of the space of tt-$(n,k)$ trades, i.e. as the irreducible decomposition of the kernel of the inclusion matrix $W_{t,k}$.
@@
Let $t<k\leq n/2$. Then vector space of tt-$(n,k)$ trades decomposes as the direct sum $\langle\mathfrak{T}_{t,k,n}\rangle\oplus\cdots\oplus\langle\mathfrak{T}_{k-1,k,n}\rangle$. [^1] In [7, Theorem 11] the result is stated without assumptions on $k,n$. However, in its proof [7, Theorem 9] is used which has such assumptions. These yield $2k-1\leq n$.
@@
From Theorem 3.4 we have $S^{(n-i-1,i+1)}\cong\langle\mathfrak{T}_{i,k,n}\rangle$ for all $i=t,\dots,k-1$. (The hypothesis $k\leq n/2$ is needed in order to apply Theorem 3.4 for $t$ replaced by k1k-1. Also it is needed so that (nk,k)(n-k,k) is a partition.)∎
For a tt-$(n,k)$ trade $T$ we denote by $\Pi(T)$ the orbit of $T$ under the action of the symmetric group $\mathfrak{S}_{n}$. The following decomposition result was proved in [7, Theorem 2].
@@
where $I\subseteq\{t,\dots,k-1\}$ consists of integers $i$ such that there is an $i$-trade in $\langle\Pi(T)\rangle$ which is not an $(i+1)$-trade. In particular, $\langle\Pi(T)\rangle$ is the whole vector space of tt-$(n,k)$ trades if and only if $I=\{t,\dots,k-1\}$.
@@
From Corollary 3.6 it follows that the irreducible decomposition of the $\mathfrak{S}_{n}$-module $\langle\Pi(T)\rangle$ is of the form $\bigoplus_{i\in I}\langle\mathfrak{T}_{i,k,n}\rangle$, where $I$ is a nonempty subset of $\{t,\dots,k-1\}.$ For every $i\in I$, the set $\mathfrak{T}_{i,k,n}$ consists of ii-$(n,k)$ trades. These are not (i+1)(i+1)-$(n,k)$ trades; indeed, by Theorem 3.4 we have $\langle\mathfrak{T}_{i,k,n}\rangle=S^{(n-i-1,i+1)}$ and this Specht module is not contained in the kernel of the map $\psi_{k}^{(k-i-1)}:M_{k}\to M_{i+1}$ according to (3.3) for $i+1$ in place of tt.∎
A classical result of Granert and Jurkat states that the $\mathbb{Z}$-module spanned by all tt-$(n,k)$ trades is spanned by the permutations of any nonzero minimal tt-$(n,k)$ trade. As an immediate corollary of Theorem 3.4 we recover the vector space version of this result.
@@
Let $t<k\leq n/2$. The vector space of tt-$(n,k)$ trades is spanned by the permutations of any nonzero minimal tt-$(n,k)$ trade.
@@
Since every total trade is a sum of minimal tt-$(n,k)$ trades, we see from Corollary 3.6 that the vector space of tt-$(n,k)$ trades is spanned by the minimal tt-$(n,k)$ trades. From this and the observation that the symmetric group $\mathfrak{S}_{n}$ acts transitively on minimal trades the result follows. ∎